OpenStudy (anonymous):

Sharmi ran 3 miles in 27 minutes. How long would it take her to run 5 miles? A. 16 minutes B. 29 minutes C. 36 minutes D. 45 minutes

OpenStudy (anonymous):

3 times what equals 27?

OpenStudy (anonymous):

just do that and multiply ur answer by 5.

OpenStudy (anonymous):

the answer is 45.

OpenStudy (anonymous):

hope i helped!

OpenStudy (anonymous):

\[\frac{27}{3}=\frac{x}{5}\] \[9=\frac{x}{5}\] \[x=9\times 5\]